What the unit for rate constants if overall Reaction order =0,1,2 and 3 ?
Answer #1

Zero order = M/s or mol.lit-1/s

First order = sec-1

Second order= M-1/s or lit.mol-1/s

Third order = lit2.Mol-2/s or M-2/s